sec (theta)=negative two over the square root of three


solve for theta



1
Expert's answer
2021-08-09T05:44:24-0400
"\sec \theta = -\frac{2}{\sqrt 3}"

Recall:


"\sec \theta = \frac{1}{\cos \theta}"

Therefore,


"\frac{1}{\cos \theta} = -\frac{2}{\sqrt 3}\\
\cos \theta = -\frac{\sqrt 3}{2}\\
\theta = \cos^{-1}\Big(-\frac{\sqrt 3}{2}\Big)\\
\theta = \frac{5 \pi}{6}"

The cosine function is negative in the second and third quadrants. To find the second solution, subtract the reference angle from "2 \pi" to find the solution in the third quadrant.



"\theta = 2 \pi - \frac{5 \pi}{6}\\
\theta= \frac{7 \pi}{6}"

The period of the "\cos\theta"  function is "2 \pi" so values will repeat every "2 \pi" radians in both directions.


"\therefore \qquad \theta= \frac{5 \pi}{6}+ 2 \pi n, \frac{7 \pi}{6}+ 2 \pi n \text{ (for any integer } n)"